OpenStudy (anonymous):

Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary. 30 mi / h = ____ ft / min a. 2640 c. 880 b. 264 d. 2112

OpenStudy (anonymous):

@austinL

OpenStudy (austinl):

Are you familiar with dimensional analysis?

OpenStudy (anonymous):

Somewhat.

OpenStudy (austinl):

Ok, we have this \[\frac{30 miles}{1 hour}=\frac{? feet}{1 minute}\] How many feet in one mile?

OpenStudy (anonymous):

5280?

OpenStudy (austinl):

Correct, so we would have \[\frac{30miles}{1 hour}\times\frac{1mile}{5280feet}...\] Now we need to get the hours to minutes. How many minutes in one hour?

OpenStudy (anonymous):

60

OpenStudy (austinl):

Correct, so now we have \[\frac{30 miles}{1hour}\times\frac{5280feet}{1mile}\times\frac{1hour}{60minutes}=?\] I made an error before it is 5280/1

OpenStudy (anonymous):

Okay? So, now what?

OpenStudy (austinl):

Multiply the numbers across the top, and multiply the numbers across the bottom. \[\frac{30\times5280\times1}{1\times1\times60}=?\]

OpenStudy (anonymous):

158,400/60?

OpenStudy (austinl):

Can that be simplified?

OpenStudy (anonymous):

1504/.6?

OpenStudy (austinl):

\[\frac{158,400}{60}=2640ft/min\]

OpenStudy (anonymous):

Thank you. (:

OpenStudy (austinl):

No problem!